### v2

# Memory-adaptive Depth-wise Heterogenous Federated Learning

Abstract
Federated learning is a promising paradigm that allows multiple clients to collaboratively train a model without sharing the local data. However, the presence of heterogeneous devices in federated learning, such as mobile phones and IoT devices with varying memory capabilities, would limit the scale and hence the performance of the model could be trained. The mainstream approaches to address memory limitations focus on width-slimming techniques, where different clients train subnetworks with reduced widths locally and then the server aggregates the subnetworks. The global model produced from these methods suffers from performance degradation due to the negative impact of the actions taken to handle the varying subnetwork widths in the aggregation phase. In this paper, we introduce a memory-adaptive depth-wise learning solution in FL called FeDepth, which adaptively decomposes the full model into blocks according to the memory budgets of each client and trains blocks sequentially to obtain a full inference model. Our method outperforms state-of-the-art approaches, achieving 5% and more than 10% improvements in top-1 accuracy on CIFAR-10 and CIFAR-100, respectively. We also demonstrate the effectiveness of depth-wise fine-tuning on ViT. Our findings highlight the importance of memory-aware techniques for federated learning with heterogeneous devices and the success of depth-wise training strategy in improving the global model’s performance.
Introduction
Federated Learning (FL) is a popular distributed learning paradigm that can address decentralized data and privacy-preserving challenges by collaboratively training a model among multiple local clients without centralizing their private data (McMahan et al. 2017; Kairouz et al. 2021). FL has gained widespread interest and has been applied in numerous applications, such as healthcare (Du Terrail et al. 2022), anomaly detection (Zhang et al. 2021), recommendation system (Lin et al. 2020b), and knowledge graph completion (Zhang et al. 2022). However, a defining trait of FL is the presence of heterogeneity — 1) data heterogeneity, where each client may hold data according to a distinct distribution, leading to a sharp drop in accuracy of FL (Zhao et al. 2018), and 2) heterogeneous clients, which are equipped with a wide range of computation and communication capabilities — challenges the underlying assumption of conventional FL setting that local models have to share the same architecture as the global model (Diao, Ding, and Tarokh 2021). In the last five years, data heterogeneity has been largely explored in many studies (Karimireddy et al. 2019; Lin et al. 2020a; Li et al. 2020a; Seo et al. 2020; Acar et al. 2021; Zhu, Hong, and Zhou 2021; Li, He, and Song 2021; Tan et al. 2022). However, only a few works aim to address the problem of heterogeneous clients, particularly memory heterogeneity in FL  (Diao, Ding, and Tarokh 2021; Hong et al. 2022).
One solution to heterogeneous clients is to use the smallest model that all clients can train, but this can severely impact FL performance as larger models tend to perform better  (Frankle and Carbin 2019; Neyshabur et al. 2019; Bubeck and Sellke 2021). Another approach is to prune channels of the global model for each client based on their memory budgets and average the resulting local models to produce a full-size global model  (Diao, Ding, and Tarokh 2021; Hong et al. 2022; Horvath et al. 2021). However, such approaches suffer from the issue of under-expression of small-size models, since the reduction in the width of local models can significantly degrade their performance due to fewer parameters  (Frankle and Carbin 2019). The negative impact of aggregating small-size models in FL is also verified by our case studies in Section Behaviors of Sub-networks in Prior Works.
Considering the exceptional performance of the full-size model, we aim to provide an algorithmic solution to enable each client to train the same full-size model and acquire adequate global information in FL. Specifically, we propose memory-adaptive depth-wise learning, where each client sequentially trains blocks of a neural network based on the local memory budget until the full-size model is updated. To ensure the classifier layer’s supervised signal can be utilized for training each block, we propose two learning strategies: 1) incorporating a skip connection between training blocks and the classifier, and 2) introducing auxiliary classifiers. Our method is suitable for memory-constrained settings as it does not require storing the full intermediate activation and computing full intermediate gradients. Additionally, it can be seamlessly integrated with most FL algorithms, e.g., FedAvg (McMahan et al. 2017) and FedProx (Li et al. 2020a).
Apart from providing adaptive strategies for low-memory local training, we investigate the potential of mutual knowledge distillation (Hinton et al. 2015; Zhang et al. 2018) to address on-the-fly device upgrades or participation of new devices with increased memory capacity. Lastly, we consider devices with extremely limited memory budgets such that some blocks resulting from the finest network decomposition cannot be trained. We propose a partial training strategy, where some blocks that are close to the input sides are never trained throughout. The main contributions of our work are summarized as follows.
  1. 1.
    Through comprehensive analysis of memory consumption, we develop two memory-efficient training paradigms that empower each client to train a full-size model for improving the global model’s performance.
  2. 2.
    Our framework is model- and optimizer-agnostic. The flexibility allows for deployment in real-world cross-device applications, accommodating clients with varying memory budgets on the fly.
  3. 3.
    Our proposed approach is not sensitive to client participation resulting from unstable communication because we learn a unified model instead of different local models as in prior works.
  4. 4.
    Experimental results demonstrate that the performance of the proposed methods is better than other FL baselines regarding top-1 accuracy in scenarios with heterogeneous memory constraints and diverse non-IID data distributions. We also show the negative impact of sub-networks using width-slimming techniques.
Related Work
Federated Learning
FL emerges as an important paradigm for learning jointly among clients’ decentralized data (Konečnỳ et al. 2016; McMahan et al. 2017; Li et al. 2020a; Kairouz et al. 2021; Wang et al. 2021a). One major motivation for FL is to protect users’ data privacy, where users’ raw data are never disclosed to the server and any other participating users (Abadi et al. 2016; Bonawitz et al. 2017; Sun et al. 2019). Partly opened by the federated averaging (FedAvg) (McMahan et al. 2017), a line of work tackles FL as a distributed optimization problem where the global objective is defined by a weighted combination of clients’ local objectives (Mohri, Sivek, and Suresh 2019; Li et al. 2020a; Reddi et al. 2020; Wang et al. 2020b). The federated learning paradigm of FedAvg has been extended and modified to support different global model aggregation methods and different local optimization objectives and optimizers (Yurochkin et al. 2019; Reddi et al. 2020; Wang et al. 2021a, b, 2020a). Theoretical analysis has been conducted, which demonstrated that federated optimization enjoys convergence guarantees under certain assumptions (Li et al. 2020b; Wang et al. 2021a).
Device Heterogeneity in Federated Learning.
Especially for cross-device FL, it is a natural setting that client devices are with heterogeneous computation power, communication bandwidth, and/or memory capacity. A few research efforts have been paid to designing memory heterogeneity-aware FL algorithms. HeteroFL (Diao, Ding, and Tarokh 2021) and FjORD (Horvath et al. 2021) allows model architecture heterogeneity among participating clients via varying model widths. The method bears similarity to previously proposed slimmable neural network  (Yu et al. 2018; Yu and Huang 2019) where sub-networks with various widths and shared weights are jointly trained with self-distillation (Zhang, Bao, and Ma 2021). SplitMix (Hong et al. 2022) tackles the same device heterogeneity problem via learning a set of base sub-networks of different sizes among clients based on their hardware capacities, which are later aggregated on-demand according to inference requirements. While recent studies like InclusiveFL (Liu et al. 2022) and DepthFL (Kim et al. 2023) have also embraced a layer-wise training approach, they allocate model sizes to clients primarily based on a fixed network depth, e.g., taking 2 layers as a computation block. This method does not accurately represent on-device capabilities during network splitting, as layers at varying depths have distinct computation and memory costs.
In summary, prior research efforts mainly focus on partitioning the global model weights among participating users given their hardware resource constraints, which may be imprecisely evaluated. Consequently, each local user only accesses a portion of the global model. In this work, however, we seek a holistic approach to handling device heterogeneity in FL without the need for partitioning model weights. Instead, our approach allows each device to train a full-size model in a sequential manner.
Empirical Study
Preliminaries
This section briefly reviews prior significant and open-sourced works that aim to address heterogeneous clients in FL, including HeteroFL (Diao, Ding, and Tarokh 2021) and SplitMix (Hong et al. 2022). We then conduct an extensive analysis of memory consumption of training a neural network, which has not been explored thoroughly in the FL community.
Width-scaling FL for heterogeneous clients.
Existing works such as HeteroFL (Diao, Ding, and Tarokh 2021) and SplitMix (Hong et al. 2022) address memory heterogeneity by pruning a single global model in terms of channels, creating heterogeneous local models. HeteroFL is the first work in FL that tackles memory heterogeneity via the width-scaling approach but still produces a full-size global model. However, HeteroFL suffers from two major limitations: 1) partial model parameters are under-trained because only partial clients and data are accessible for training the full-size model; 2) small models’ information tends to be ignored because of their small-scale parameters. SplitMix was then proposed to address these two issues, which first splits a wide neural network into several base sub-networks for increasing accessible training data, then boosts accuracy by mixing base sub-networks.
Memory consumption analysis.
Training a neural network with backpropagation consists of feedforward and backward passes (Rumelhart, Hinton, and Williams 1986). A feed-forward pass over each block of a neural network generates an activation or output. These intermediate activations are stored in the memory for the backward pass to update the neural network. Although several works of literature (Sohoni et al. 2019; Gomez et al. 2017; Raihan and Aamodt 2020; Chen et al. 2021) demonstrate that activations usually consume most of the memory in standard training of a neural network as shown in Figure 1, HeteroFL and SplitMix merely consider the number of model parameters as the memory budget in their experiments. Specifically, they divide clients into groups that are capable of different widths, e.g., a $\frac{1}{8}$-width neural network, which costs approximately $\frac{1}{8}$ activations but only around $\frac{1}{8^{2}}$ model parameters compared to the full-size neural network.
Figure 1: Training memory consumption for left: WideResNet on CIFAR-10 and right: DC-Transformer on IWSLT’14 German to English. Data source: (Sohoni et al. 2019).
Behaviors of Sub-networks in Prior Works
In this section, we analyze the behaviors and influences of sub-networks in HeteroFL and SplitMix with respect to the performance of the global model.
Experimental setting.
We follow the configuration on the CIFAR-10 dataset in SplitMix (Hong et al. 2022), where 10 out of 100 clients participate in training in any given communication round, and each client has three classes of the data. In our case studies, we divide clients into four groups with $\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1\}$-width sub-networks in HeteroFL, and into two groups with $\{r,1\}$ in SplitMix, where $r=\{\frac{1}{16},\frac{1}{8},\frac{1}{4},\frac{1}{2}\}$. Our observations are summarized below.
  1. 1.
    Small sub-networks make negative contributions in HeteroFL. Figure 2 (left) presents typical examples of HeteroFL (Diao, Ding, and Tarokh 2021) under non-IID settings. The orange line represents the default setting of HeteroFL, where all sub-networks of different widths are aggregated. The other lines indicate specific size of sub-networks that are not aggregated. For example, the green line indicates that the smallest ($\frac{1}{8}$-width) sub-networks do not participate in aggregation. We observe that the global model obtained via aggregating small sub-networks consistently has worse performance than the global model obtained via only aggregating the full-size neural networks, indicating that small size sub-networks make negative contributions.
  2. 2.
    Small sub-networks limit global performance in SplitMix. Figure 2 (right) depicts the prediction performance of the global model in SplitMix (Hong et al. 2022) by mixing base neural networks with different-width. It clearly illustrated that slimmer base neural networks produce a less powerful global model. Intuitive reasoning is that combining very weak learners leads to an ensemble model with worse generalization.
  3. 3.
    The full-size net makes a difference. Inadequate presence of the full-size models incurs degradation of validation accuracy as shown in Figure 2. Besides, in real-world FL systems, communication can be unstable, and clients with the largest memory budgets may not be available in each round of communication (Bonawitz et al. 2017). This constraint limits the practicality of both HeteroFL and SplitMix.
Figure 2: Performance of the global model in HeteroFL (left) and SplitMix (right) with the varying-width base model, respectively.
Methodology
Inspired by the observation in the previous section, we introduce a memory-efficient framework FeDepth to train full-size neural networks with memory budget constraints in the FL setting. FeDepth aims to empirically solve the optimization problem $\min_{{\cal W}}F({\cal W}):=\sum_{k=1}^{K}p_{k}F_{k}({\cal W})$. Here, $F_{k}$ represents the loss function on the $k$th clients. $p_{k}>0$ for all $k$ and $\sum_{k=1}^{K}p_{k}=1$. FeDepth features memory-adaptive decomposition, where a neural network is decomposed into blocks based on the memory consumption and local clients’ memory budgets. An implict assumption made is that all blocks can be trained locally after the decomposition. To further address extreme case that some blocks still cannot fit into the memeory even after the finnest decomposition, FeDepth integrates the partial training strategy into the local training stage. We also consider the possibilty that some clients with rich memory budgets may suffer from memory underutilization, hence a variant of FeDepth is propsed to use mutual knowledge distillation (Zhang et al. 2018) to boost the performance and fully exploit the local memory of clients.
FeDepth and Its Variants
Algorithm 1 FeDepth
1:Total number of clients $K$; participation rate $\gamma$; number of communication rounds $R$.
2:Model parameter ${\cal W}^{0}$.
3:for $t=0,\dots,R-1$ communication rounds do
4:     Sample a subset ${\cal S}^{t}$ of clients with $|{\cal S}^{t}|=\lceil\gamma K\rceil$.
5:     Broadcast ${\cal W}^{t}$ to clients $k\in{\cal S}^{t}$.
6:     for each client $k\in{\cal S}^{t}$ in parallel do
7:         ${\cal W}^{t+1}_{k}\leftarrow\text{ClientUpdate}({\cal W}^{t},k)$.
8:     end for
9:     Aggregate as ${\cal W}^{t+1}=\sum_{k\in{\cal S}^{t}}\frac{p_{k}}{\sum_{k^{\prime}\in{\cal S}^{t}}p_{k^{\prime}}}{\cal W}^{t+1}_{k}.$
10:end for
11:procedure ClientUpdate(${\cal W}^{t},k$)
12:     for $j=1,\cdots,J_{k}$ do
13:         Approximately solve the problem (1).
14:     end for
15:     Set $\phi_{k}^{t+1}=\phi_{J}^{t+1}$
16:     Return ${\cal W}^{t}_{k}=\{\theta^{t+1}_{k,1},\cdots,\theta^{t+1}_{k,J_{k}},\phi_{k}^{t+1}\}$.
17:end procedure
Memory-adaptive network decomposition.
Since various clients could have drastically different memory budgets, FeDepth conducts local training in a memory-adaptive manner. Specifically, for the $k$-th client the full model ${\cal W}$ is decomposed to into $J_{k}+1$ blocks, i.e., ${\cal W}=\{\theta_{k,1},\cdots,\theta_{k,J_{k}},\phi\}$, where $\{\theta_{k,j}\}_{j=1}^{J_{k}}$ and $\phi$ denote body and head of the neural network, respectively. Note that $\theta_{k,j}$ can be different from $\theta_{k^{\prime},j}$ for any $(k,k^{\prime},j)$ triple, and the number of parameters contained in $\theta_{k,j}$ is solely determined by the $k$th client’s memory budget, hence FeDepth is memory-adaptive. In practice, the model decomposition can be determined for each client before training via the estimating memory consumption (Gao et al. 2020). See Figure 3 for an illustration. Suppose the full-size model is composed of 6 layers, where each of layer costs memory of $\{3,2,1,0.5,0.5,0.5\}$ GB, respectively. Assume the $k$th and $k^{\prime}$th client has 3 GB and 5 GB memory budget, respectively. Then, client $k$ has $J_{k}=3$ and client $k^{\prime}$ has $J_{k^{\prime}}=2$ trainable blocks, respectively. That is, client $k^{\prime}$ will start with training the first two blocks, then the remaining four blocks.
Depth-wise sequential learning.
Once the decomposition is determined, the $k$-th client at the $t$-th round, locally solves $J_{k}$ subproblems in a block-wise fashion, i.e., for all $j\in\{1,\dots,J_{k}\}$,
$(\theta^{t+1}_{k,j},\phi_{j}^{t+1})\in\arg\min_{\theta_{k,j},\phi}{\cal L}(\theta_{k,j},\phi;\{z_{j-1,i}^{t+1},y_{i}\}_{i=1}^{n_{k}}),$ (1)
where ${\cal L}$ is a loss function, e.g, cross-entropy; $\{z_{j-1,i}^{t+1}\}_{i=1}^{n_{k}}$ are activations obtained after the local training samples forward-passed through the first $j$ blocks, i.e., $z_{j-1,i}^{t+1}=f(x_{i};\{\theta_{k,\ell}^{t+1}\}_{\ell=1}^{j-1})$ for $i\in[n_{k}]$ and $f(\cdot;\{\theta_{k,\ell}^{t+1}\}_{\ell=1}^{j-1})$ is the neural network up to the first $j-1$ blocks; $n_{k}$ is the total number of training samples. Specifically, $z_{0,i}^{t+1}=x_{i}$ for all $i\in[n_{k}]$. Problem (1) can be solved by any stochastic gradient-based method. When solving for the $j$-th subproblem at the $t$th round, to fully leverage the global model ${\cal W}^{t}$’s body and the locally newly updated head, we use $(\theta_{j}^{t},\phi_{j-1}^{t+1})$ as the initial point. See the data flow in Figure 4 when performing the training for the $j$th block. We remark that when the memory budget permits, the activation $\{z_{j-1,i}^{t+1}\}_{i=1}^{n_{k}}$, which no longer requires the gradient, could be buffered to compute $\{z_{j,i}^{t+1}\}_{i=1}^{n_{k}}$ once the $\theta_{k,j}^{t+1}$ is obtained, hence saving the redundant forward-pass from the first block to the $j$th block. Also, the number computation required for approximately solving $J_{k}$ subproblems in the form of (1) should be similar to approximately solving the problem $\min_{{\cal W}}F_{k}({\cal W})$ if we perform the same number of local updates, and ignore the negligible computation overhead in updating in the head $\phi$. This is because the amount of computation required by one gradient evaluation $\nabla_{{\cal W}}F_{k}$ is equivalent to that of the summation of gradient evaluation of $\{\nabla_{\theta_{j}}{\cal L}\}_{j=1}^{J_{k}}$.
Figure 3: Memory-adaptive neural network decomposition. The second row represents the full neural network with the block size indicating the memory consumption. The first and third rows explain the neural network decomposition based on different clients’ memory budgets.
Figure 4: An example of depth-wise sequential learning. There are three training steps: 1) training the first block and the classifier with the skip connection (He et al. 2016a); 2) freezing the updated first block and using its activation to train the second block and the classifier with the skip connection; 3) freezing the updated first two blocks and using the activation of the second block to train the third block and the classifier.
Memory-efficient Inference.
Depth-wise inference follows the similar logic of the frozen-then-pass forward in depth-wise training. Specifically, for each input $x$, we store the activation $z_{j}$ in the hard drive and discard the predecessor activation $z_{j-1}$. Then we can reload $z_{j}$ into memory as the input and get the activation $z_{j+1}$. The procedure is repeated until the prediction $\hat{y}$ is obtained.
We end this section by giving the detailed algorithmic description in Algorihtm 1.
Handle Extrem Memory Constraints with Partial Tranining
According to the memory consumption analysis in the Empirical Study, the memory bottleneck of training a neural network is related to the block with the largest activations, which some devices may still not afford. These “large” blocks are usually the layers close to the input side. To tackle this issue, we borrow the idea of partial training in FeDepth, where we skip the first few blocks that are too large to fit even after the finnest blocks decomposition. This mechanism would not incur significant performance degradation because the input-side layers learn similar representations on different datasets (Kornblith et al. 2019), and clients with sufficient memory will provide model parameters of these input-side layers in the FL aggregation phase. Figure 5 emeprically validates such a strategy. We train a customized 14-layer ResNet (13 convolution layers and one classifier) on MNIST with 20 clients under non-IID distribution, respectively and measure the similarity of neural network representations, using both canonical correlation analysis (CCA) and centered kernel alignment (CKA) (Kornblith et al. 2019).
Figure 5: Correspondences between layers of different local neural networks trained from private datasets in FL under non-IID distribution. We observe that early layers, but not later layers, learn similar representations.
Exploit Sufficient Memory with Mutual Knowledge Distillation
Previous works on heterogeneous FL ignore the situation where some clients with rich computing resources may participate in the federated training on the fly. Their sufficient memory budget could be potentially utilized to improve the performance of FL via regularizing local model training (Mendieta et al. 2022; Li et al. 2020a). Ensembling multiple models is an effective way to improve generalization and reduce variance (Shi et al. 2021; Nam et al. 2021). However, considering each model is independently trained in ensemble learning methods, we have to upload/download all of these models in FL settings leading to a significant communication burden. Therefore, we design a new training and aggregation method based on mutual knowledge distillation (MKD) (Hinton et al. 2015; Zhang et al. 2018), where all student neural networks learn collaboratively and teach each other. Therefore, clients with sufficient memory only need to upload one of the local models to the server for aggregation because the knowledge consensus achieved among all models through distillation. Formally, assume the $k$th client has a rich memory budget to train $M>1$ models. Then locally it solves
$\min_{\{{\cal W}_{k}^{1},\cdots,{\cal W}_{k}^{M}\}}$ $\frac{1}{M}\sum_{m=1}^{M}F_{k}({\cal W}_{k}^{m})+\frac{1}{M-1}\sum_{m^{\prime}\neq m}^{M}\operatorname{\boldsymbol{\mathbf{KL}}}\left(\mathbf{h}^{m^{\prime}}\|\mathbf{h}^{m}\right),$
where $\mathbf{h}^{m}$ are logits calculated from the model ${\cal W}_{m}$ over the local training set and $\operatorname{\boldsymbol{\mathbf{KL}}}$ is the Kullback Leibler Divergence. More concretely, $\operatorname{\boldsymbol{\mathbf{KL}}}\left(\mathbf{h}^{m^{\prime}}\|\mathbf{h}^{m}\right)=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\operatorname{\boldsymbol{\mathbf{KL}}}(h_{i}^{m^{\prime}}\|h_{i}^{m})$, where $h_{i}^{m}$ is the logits of the $i$th sample computed over model ${\cal W}_{m}$.
Experiments
Depth Memory Width Memory
[…]
Experimental Setups
Datasets and data partition. Our experiments are mainly conducted on CIFAR-10 and CIFAR100 (LeCun et al. 1998; Krizhevsky and Hinton 2009). Extensive results are attached in the appendix. To simulate the non-IID setting with class imbalance, we follow (Yurochkin et al. 2019; Acar et al. 2021; Gao et al. 2022) to distribute each class to clients using the Dirichlet distribution with $\alpha(\lambda)$, where $\lambda=\{0.3,1.0\}$. Besides, we adopt pathological non-IID data partition $\beta(\Lambda)$ that is used by the selected baselines – HeteroFL and SplitMix (Diao, Ding, and Tarokh 2021; Hong et al. 2022), where each device has unique $\Lambda$ labels with $\Lambda=\{2,5\}$ for CIFAR-10 while $\Lambda=\{10,30\}$ for CIFAR-100. We note that the balanced data partition is applied by default, which makes each client holds the same number of examples. The unbalanced $\alpha_{u}(\lambda)$ non-IID, where clients may have a different amount of samples with different feature distribution skew, is also used to evaluate the stability of FeDepth.
[…]
  • (Fair). The memory budgets depend on the hidden channel shrinkage ratio, $r=\{\frac{1}{6},\frac{1}{3},\frac{1}{2},1\}$, and are uniformly distributed into clients. It means 1/4 of clients can train a PreResNet-20 model within a maximal width of $\frac{1}{6}$-, $\frac{1}{3}$-, $\frac{1}{2}$- and full width, respectively.
  • (Lack). $r=\{\frac{1}{8},\frac{1}{6},\frac{1}{2},1\}$. 1/4 of clients with limited memory adopt a partial training strategy.
  • (Surplus). $r=\{\frac{1}{6},\frac{1}{3},\frac{1}{2},2\}$. 1/4 of clients with sufficient memory can apply MKD.
Table 2: Test results (top-1 accuracy) under balanced non-IID data partitions using PreResNet-20. Grey texts indicate that the training cannot conform to the pre-defined budget constraint. If not specified, FedAvg denotes the results with ×$\times\min(r)$width network. We highlight the best results with Blue Shadow, Red Shadow, and Bold in the scenarios including clients equipped with fairly sufficient, insufficient and abundant memory, respectively.
Budget Method CIFAR-10 CIFAR-100
$\alpha(0.3)$ $\alpha(1.0)$ $\beta(2)$ $\beta(5)$ $\alpha(0.3)$ $\alpha(1.0)$ $\beta(10)$ $\beta(30)$
Unrealistic FedAvg (×$\times 1$) 63.44 $\pm$ 2.95 69.77 $\pm$ 1.48 34.10 $\pm$ 1.68 68.33 $\pm$ 0.75 28.14 $\pm$ 0.48 31.28 $\pm$ 0.17 31.56 $\pm$ 0.28 43.54 $\pm$ 0.16
Fair FedAvg (×$\times\frac{1}{6}$) 47.66 $\pm$ 2.67 53.10 $\pm$ 1.27 25.82 $\pm$ 1.34 54.02 $\pm$ 0.96 15.83 $\pm$ 0.25 16.39 $\pm$ 0.08 13.89 $\pm$ 0.21 17.96 $\pm$ 0.08
HeteroFL 56.35 $\pm$ 3.44 60.48 $\pm$ 3.23 26.47 $\pm$ 2.98 57.51 $\pm$ 2.76 19.20 $\pm$ 3.35 23.62 $\pm$ 3.48 20.52 $\pm$ 2.14 32.04 $\pm$ 2.56
SplitMix 55.63 $\pm$ 2.60 58.70 $\pm$ 1.33) 28.70 $\pm$ 1.92 58.15 $\pm$ 2.03 22.61 $\pm$ 0.88 24.86 $\pm$ 0.44 19.55 $\pm$ 0.79 26.89 $\pm$ 0.32
DepthFL 58.52 $\pm$ 1.76 58.55 $\pm$ 1.27 29.21 $\pm$ 1.73 59.29 $\pm$ 1.60 23.56 $\pm$ 1.45 25.27 $\pm$ 0.94 23.14 $\pm$ 1.26 33.46 $\pm$ 1.08
FeDepth 60.62 $\pm$ 1.84 64.95 $\pm$ 1.56 30.32 $\pm$ 2.19 61.26 $\pm$ 1.83 26.20 $\pm$ 1.61 27.57 $\pm$ 1.67 24.47 $\pm$ 1.42 36.35 $\pm$ 1.21
$m$-FeDepth 60.03 $\pm$ 1.90 65.49 $\pm$ 2.18 32.50 $\pm$ 2.42 64.18 $\pm$ 1.95 30.53 $\pm$ 1.30 32.85 $\pm$ 1.85 25.15 $\pm$ 1.58 38.36 $\pm$ 1.34
[…]
HeteroFL 54.05 $\pm$ 3.72 58.03 $\pm$ 3.58 25.83 $\pm$ 3.35 55.68 $\pm$ 3.33 18.40 $\pm$ 3.45 21.79 $\pm$ 3.52 18.70 $\pm$ 1.34 29.33 $\pm$ 2.12
SplitMix 49.18 $\pm$ 2.88 52.95 $\pm$ 1.60 24.47 $\pm$ 2.29 52.40 $\pm$ 1.48 21.20 $\pm$ 1.23 22.60 $\pm$ 1.01 17.71 $\pm$ 1.15 22.58 $\pm$ 0.64
DepthFL 56.04 $\pm$ 2.03 57.95 $\pm$ 1.50 27.02 $\pm$ 2.12 57.13 $\pm$ 2.10 22.57 $\pm$ 1.68 24.18 $\pm$ 1.12 21.20 $\pm$ 1.33 31.14 $\pm$ 1.45
FeDepth 58.63 $\pm$ 2.11 62.83 $\pm$ 1.83 29.01 $\pm$ 2.54 59.76 $\pm$ 1.28 25.31 $\pm$ 2.52 26.97 $\pm$ 1.34 22.95 $\pm$ 1.29 34.71 $\pm$ 1.78
$m$-FeDepth 57.56 $\pm$ 2.18 62.61 $\pm$ 2.45 28.34 $\pm$ 2.79 62.73 $\pm$ 2.44 30.39 $\pm$ 1.68 31.08 $\pm$ 1.57 23.58 $\pm$ 1.41 37.34 $\pm$ 1.89
[…]
$m$-FeDepth 62.30 $\pm$ 1.25 66.65 $\pm$ 1.65 33.82 $\pm$ 1.85 67.57 $\pm$ 1.89 32.55 $\pm$ 1.13 37.00 $\pm$ 1.52 29.20 $\pm$ 1.70 40.42 $\pm$ 1.58
Implementation and evaluation. We compare FeDepth and its variants with several methods, including FedAvg (McMahan et al. 2017), HeteroFL (Diao, Ding, and Tarokh 2021), SplitMix (Hong et al. 2022) and DepthFL  [^2] We reproduced this algorithm to conform to our predefined memory budgets, rather than the original fixed-depth allocation. (Kim et al. 2023) in terms of the average Top-1 Accuracy over 5 different runs. The memory budgets are uniformly distributed to 100 clients. All experiments perform 500 communication rounds with a learning rate of 0.1, local epochs of 10, batch size of 128, SGD optimizer, and a cosine scheduler.
Table 2: Test results (top-1 accuracy) under balanced non-IID data partitions using PreResNet-20. Grey texts indicate that the training cannot conform to the pre-defined budget constraint. If not specified, FedAvg denotes the results with ×$\times\min(r)$width network. We highlight the best results with Blue Shadow, Red Shadow, and Bold in the scenarios including clients equipped with fairly sufficient, insufficient and abundant memory, respectively.